Operation Summary for December 11, 2009 - December 14, 2009 |
The Integrated Luminosity for the period from 12/7/09 to 12/14/09 was 36.72 pb-1. Friday Dec. 11 The day shift began with Operations monitoring the Tevatron (TeV) store 7433 (208E30) and the Antiproton Source stack (103.4E10), with no Recycler stash due to a Pelletron access, with the Linac operating normally, with Booster and Main Injector (MI) off due to accesses, with the NuMI and MiniBooNE experiments off due to the MI and Booster accesses, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day and evening shifts, Operations reported that the Meson corrector magnet (MT5E) replacement had begun. Operations reported accesses to MI, Booster, MiniBooNE, and the MI-8 transfer line. The Booster access ended at 8:30 PM and beam was established to the Booster dump by 11:15 PM. The MI-8 line vacuum recovery was slow after the multiwire installation in the MI-8 line. The MI RF cavity (MIRF13) replacement developed an LCW leak. Saturday Dec. 12 The midnight shift began with Operations monitoring store 7433 (54.85E30) and the stack (101.3E10), with Pelletron and MI work continuing. During the midnight, day, and evening shifts, Operations reported that technicians had completed the MIRF13 cavity replacement by 2:30 AM. A Linac RF station (LRF5) tripped off; Operations called in Linac personnel who found a noisy power supply and a LCW leak, which they repaired. By noon, the vacuum had recovered in the MI-8 line. The MI-60 vacuum recovery continued. The MI access was completed by 5:45 PM. MiniBooNE resumed taking beam. Sunday Dec. 13 The midnight shift began with Operations monitoring store 7433 and the stack, and with MiniBooNE taking beam. During the midnight shift, at 4:11 AM, the MI abort kicker power supplies (KPS4A and KPS4B) were problematic. Operators reduced the reservoir voltage on KPS4B and this resolved the problem. An AP-1 quadrupole magnet (Q106) would not turn on; Pbar personnel came in and replaced a faulty sensor. Once the MI-60 vacuum was good enough for beam operations, Operations established beam to NuMI at 7:20 AM. Pbar resumed stacking at 7:40 AM. Operations terminated store 7433 at 9:50 AM. Operations prepared CDF and D0 for accesses into their collision halls. Operations put the TeV into HEP shot setup at 2 PM and established store 7437 at 3:10 PM with an average initial luminosity (AIL) of 125.0E30. Power supply personnel conducted an inspection of the MT5E magnet and reported it was ready for beam. Operations sent beam to MTest from 4:30 PM to 8 PM. Operations reported that the Pelletron had developed a BPM problem on its deceleration side, which caused its autotune program to issue incorrect values. Monday Dec. 14 The midnight shift began with Operations monitoring store 7437 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:50 AM, Operations reported that a failed BPM front end (FE) had inhibited the NuMI beam. Once the FE recovered from its reboot, the trim magnet values were restored and Operations resumed sending beam at 1:15 AM. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow a NuMI access to repair a LCW leak; a Pbar aperture scan; and a Tevatron Electron Lens (TEL2) study. |
|
More Information Comments and Suggestions |
| last modified 12/14/2009 email Fermilab |
| Security, Privacy, Legal | |